An object will fall m feet in n seconds where m = 16nsquared. How long will it
take for a quarter to hit the ground if it is dropped from
the top of a tower that is 1454 feet tall?

Answer by checkley71(8403)

M=16N^2
1454=16N^2
16N^2=1454
N^2=1454/16
N^2=91
N=SQRT91
N=9.54 SECONDS.